# Beyond the Surface: Probing the Ideological Depth of Large Language Models

### v1

@@ 3.3 Political Feature Analysis
Motivated by these findings, we turned our attention towards analyzing the Sparse Autoencoder (SAE) features, hoping to find a difference in their political feature that can explain their behavioral differences while acting as liberal or conservative.
SAEs are a useful tool for identifying monosemantic features from the vectors of residual streams of a layer form a model. SAEs achieves this by projecting the vector to a high-dimensional vector space with the goal of disentangling the complex features that are otherwise superpositioned in the outputs of a single neuron [9, 5]. Recent works already show that models’ intermediate layer store vectors responsible for their ideological behavior that are highly interpretable [14]. Our goal was to see if the SAE features from these layers can be used to quantify their political knowledge and subsequently explain why their capabilities vary when playing different roles. Although there are typically thousands of features in a single SAE, not all of them activate or are equally important for a particular behavior. Recent SAE-based steering approaches select features based on their activation values (amplitude), activation frequencies [27], and their steering strength [3]. First, we conduct a quantitative comparison to measure the raw number and activation strength of activated features in each model. We then perform a qualitative evaluation to assess the thematic coherence and predictive power of these features, to confirm how they represent genuine political concepts.
Quantitative Analysis: We used the approach of Wang et al. [27] to collect and analyze the normalized mean amplitude (activation difference) and the normalized mean frequency difference of the activated features for positive and negative data from the 874 training prompts using $\Delta a=\frac{1}{N}\sum^{N}_{i}(\bar{a}_{i}^{pos}-\bar{a}_{i}^{neg})$ and $\Delta f=f^{pos}-f^{neg}$. Where, $f^{pos,j}=\frac{1}{N}\sum^{N}_{i}\mathbb{I}(|\bar{a}_{i}^{pos,j}|>0)$ and $f^{neg,j}=\frac{1}{N}\sum^{N}_{i}\mathbb{I}(|\bar{a}_{i}^{neg,j}|>0)$.
@@
Table 2: Statistical Summary of Output Score Data
We also collected all the features that activate for the 126 statements in our evaluation set by projecting the activation vector of the target layer on the SAE space, marking the token of the statement that most activates this feature, and neglecting features activated for the bos token. In total, for the gemma-2-9b-it model, we found 18458 features, whereas for the llama-3.1-8b-it, we found only 4412 activated features out of a total possible 131K features in both SAEs. On average, Gemma contained 7.3 times more features that activate for a statement.
Next, we selected features with $\Delta f>0$ or $\Delta a>0$ from both models and evaluated the quality of these features based on their ability to steer the model’s behaviour. We got 71 and 16 features from Gemma and Llama SAEs, respectively. This is done because if the features can steer the model heavily, even a small number of features can be enough to control the model. To do this we calculated the output scores of the features, using the approach proposed by Arad et al. [3], by performing interventions during a forward pass and evaluating the change in the rank and probability assigned to tokens in the target layer. The output score is calculated using Equation 1, which is the difference between the original and the interveined rank weighted probabilities of the target layer when the model is provided with a neutral sentence (we used "In my opinion,"). $r(l^{*},\mathcal{M})$ and $p(l^{*},\mathcal{M})$ denote the rank and probability of the highest-ranked token $l^{*}$, and $|V|$ is the vocab size of the model $\mathcal{M}$.
$\begin{split}P(\mathcal{M})=(1-\frac{r(l^{*},\mathcal{M})}{|V|})p(l^{*},\mathcal{M})\\ S_{out}=P(\mathcal{M}_{h\leftarrow{\phi(h)}}-P(\mathcal{M})\end{split}$ (1)
The stats of the output scores are shown in Table 2. The scores for Gemma appear to be heavily right-skewed, with most values clustered near zero. The median and even the 75th percentile (Q3) are zerozero. However, its maximum score reaches 0.88. This suggests Gemma contains a few extremely high quality features, pulling the mean (0.002407) above the median. Whereas the maximum feature score for Llama was significantly lower.
Qualitative analysis: To assess the quality of the features, we use GPT-4o-mini as an automated evaluator to analyze the coherence and predictive validity of SAE features from each model. We make sure to ask the LLM to provide explanations justifying each evaluation, a technique shown to improve alignment with human judgments [7]. First, we collected the feature descriptions of the features of the SAE with $\Delta a$ or $\Delta f$ from Neuronpedia [^1] https://www.neuronpedia.org/. Next, we use the prompts in Appendix D, to perform two distinct, complementary evaluations:
@@
Figure 12: Evaluation scores distribution.
Figure 13: Dominant themes in the feature descriptions.
Our evaluation reveals significant differences in the SAE features from llama-3.1-8b-it and gemma-29b-it in terms of quality and utility for modeling political contexts. Figure 13 provides an initial, high-level overview of feature quality based on the two key scores: classification confidence and thematic coherence. The "Confidence Score Distribution" (left) shows that the evaluator model was able to make high-confidence classifications using features from both models, with median scores around 4.5. However, Gemma’s distribution is more clustered at the upper end (scores of 4 and 5), whereas Llama’s distribution has a longer tail towards the lower scores, suggesting slightly less consistent clarity. The "Feature Coherence Score Distribution" (right) reveals a more dramatic distinction. Llama’s features are consistently rated as having low coherence, with scores tightly clustered around a median of 2. This indicates that its feature sets are frequently perceived as a disjointed mix of unrelated concepts. In contrast, Gemma’s coherence scores are not only higher on average but also show a wider distribution, with an additional concentration of scores between 3.0 and 5.0. This indicates that Gemma is capable of generating highly coherent and interpretable feature sets that represent a unified political or social theme.
The word clouds in Figure 13 offer an interesting qualitative explanation for this coherence gap by visualizing the dominant themes in the features, identified by the evaluator from the 2nd experiment. The themes from Llama’s features are dominated by general, syntax-focused, and often ambiguous terms such as "Mixed references," "technical," "safety programming," and "procedural." While it identifies domain-specific topics like "health" and "funding," they are embedded within a context of process and structure rather than political meaning. This aligns with the low coherence scores, as the features appear to describe the surface-level components of a statement rather than its core political intent. Conversely, the dominant themes from Gemma’s features are notably centered on the language of political and social analysis. Terms like "political discourse", "policy critique", "legal discourse", and "socioeconomic" dominate its word cloud.
@@
Figure 14: Classification Confusion Matrix.
@@ 4 Discussion
Our findings establish "ideological depth" not as a philosophical abstraction, but as a measurable, biological property of Large Language Models. By comparing Llama-3.1-8B-IT and Gemma-2-9B-IT, we’ve demonstrated that models are not monolithic in their political leanings. Rather, they possess internal feature sets of vastly different size and complexity. This directly explains the behavioral differences we observe: Gemma’s rich feature space (7.3x more political features) allows it to flexibly reason from different ideological perspectives, whereas Llama’s "shallow" feature space leads it to become brittle, resulting in poor steerability.
We have identified a weird and safety-relevant behavior and provide a mechanistic explanation rooted in its internal structure. The refusal appears not to be a sophisticated ethical guardrail, but a capability deficit. This has critical implications for AI alignment. An ideologically "shallow" model might seem safer because it refuses controversial prompts, but this brittleness makes it unpredictable. Its refusal is a failure of understanding, not a demonstration of principle. Conversely, a "deep" model like Gemma, while more capable of adopting undesirable personas, is also more transparent. Its reasoning is based on a richer, more interpretable set of internal features, offering a wider surface for monitoring and intervention.
To validate that the features were truly meaningful and not just statistical noise, we used a powerful automated evaluator for classifying a statement’s political topic using only the descriptions of the activated SAE features. The evaluator’s success with Gemma’s features confirms that they form a coherent and semantically rich map of the political domain. Its features are not random artifacts, but causally relevant components of the model’s political understanding. Whereas, the evaluator’s corresponding failure with Llama’s features proves they are too generic to be useful, confirming that its ideological representations are, indeed, shallow.
@@ 5 Conclusion and Future Work
We have shown that political understanding of LLMs is quantifiable and steerability and SAEs serve as powerful tools for probing their "ideological depth." This research opens several interesting pathways for future work, directly aligned with key problems in Applied Interpretability and Model Biology:
@@
    Mechanistic Analysis of Refusal Circuits: Our work suggests Llama defaults to refusal due to a lack of features. The next step is to find the circuit responsible. Can we find a specific "capability-check" mechanism that activates a general refusal vector when the model cannot fulfill an instruction? We can use attribution graphs [16] to trace the causal pathway from a conservative instruction to the final refusal token, providing a concrete example of debugging a model’s unintended behavior.
@@
    Activation Monitoring: Gemma’s rich set of political features is a perfect testbed for building advanced monitoring tools. We can train simple probes on the most salient features (e.g., "policy critique," "socioeconomic analysis") to detect, in real-time, the type of reasoning a model is employing. This could lead to systems that can flag when a model is entering a specific ideological frame, offering a more nuanced form of safety monitoring than simple output filtering.
@@
    Investigating the Mechanics of Instability: Our analysis consistently shows that models exhibit greater response variance when instructed to adopt a conservative persona, especially under argumentative pressure (Figures 4 and 7). This raises the question: What is the mechanistic origin of this instability? Future work could trace the circuit of features activated under these high-variance conditions [2]. Are fewer, less coherent features being activated? Or are conflicting features, e.g., a "conservative policy" feature and a "safety/refusal" feature, firing simultaneously?
  4. 4.
    Intervening Fine-Tuning for Controllable Depth: Can we control the emergence of ideological depth by intervening during instruction-tuning? By ablating nascent political features or amplifying them with steering vectors during the fine-tuning process, we could attempt to train models with a deliberately "deep" or "shallow" representation. This would test our understanding of how these complex features are learned and could be a powerful tool for building more value-aligned models from the start.
Ultimately, understanding the internal biology of models is a critical step towards building reliable and alignable models. This work provides a concrete methodology for moving beyond surface-level behavior and beginning to map the complex, latent structures that drive it.
Appendix A Factor Analysis of LLM Responses

### v2

@@ 3.3 Political Feature Analysis
Motivated by these findings, we turned our attention towards analyzing the Sparse Autoencoder (SAE) features, hoping to find a difference in their political feature that can explain their behavioral differences while acting as a liberal or a conservative.
SAEs achieves this by projecting the vector to a high-dimensional vector space with the goal of disentangling the complex features that are otherwise superpositioned in the outputs of a single neuron elhage2022toymodelssuperposition; bricken2023monosemanticity. Our goal was to see if the SAE features from these layers can be used to quantify their political knowledge and subsequently explain why their capabilities vary when playing different roles. Although there are typically thousands of features in a single SAE, not all of them activate or are equally important for a particular behavior. Recent SAE-based steering approaches select features based on their activation values (amplitude), activation frequencies wang2025promptengineeringrobustbehavior, and their steering strength arad2025saesgoodsteering. First, we conduct a quantitative comparison to measure the raw number and activation strength of activated features in each model. We then perform a qualitative evaluation to assess the thematic coherence and predictive power of these features, to confirm how they represent genuine political concepts.
Quantitative Analysis: We used the approach of wang2025promptengineeringrobustbehavior to collect and analyze the normalized mean amplitude (activation difference) and the normalized mean frequency difference of the activated features for positive and negative data from the 874 training prompts using $\Delta a=\frac{1}{N}\sum^{N}_{i}(\bar{a}_{i}^{pos}-\bar{a}_{i}^{neg})$ and $\Delta f=f^{pos}-f^{neg}$. Where, $f^{pos,j}=\frac{1}{N}\sum^{N}_{i}\mathbb{I}(|\bar{a}_{i}^{pos,j}|>0)$ and $f^{neg,j}=\frac{1}{N}\sum^{N}_{i}\mathbb{I}(|\bar{a}_{i}^{neg,j}|>0)$.
@@
Table 2: Summary statistics of SAE feature output scores. Gemma’s features consist of a few extremely high-quality features that pull the mean above the median despite most values being clustered near zero.
[…]
Next, we selected features with $\Delta f>0$ or $\Delta a>0$ from both models and evaluated the quality of these features based on their ability to steer the model’s behaviour. We got 71 and 16 features from Gemma and Llama SAEs, respectively. This is done because if the features can steer the model heavily, even a small number of features can be enough to control the model. To do this, we calculated the output scores of the features, using the approach proposed by arad2025saesgoodsteering, by performing interventions during a forward pass and evaluating the change in the rank and probability assigned to tokens in the target layer. The output score is calculated using Equation 1, which is the difference between the original and the interveined rank weighted probabilities of the target layer when the model is provided with a neutral sentence (we used "In my opinion,"). $r(l^{*},\mathcal{M})$ and $p(l^{*},\mathcal{M})$ denote the rank and probability of the highest-ranked token $l^{*}$, and $|V|$ is the vocab size of the model $\mathcal{M}$.
$\begin{split}P(\mathcal{M})=(1-\frac{r(l^{*},\mathcal{M})}{|V|})p(l^{*},\mathcal{M}),\\ S_{out}=P(\mathcal{M}_{h\leftarrow{\phi(h)}}-P(\mathcal{M}).\end{split}$ (1)
The stats of the output scores are shown in Table 2. The scores for Gemma appear to be heavily right-skewed, with most values clustered near zero. The median and even the 75th percentile (Q3) are zerozero. However, its maximum score reaches 0.88. This suggests Gemma contains a few extremely high quality features, pulling the mean (0.002407) above the median. In contrast, the maximum feature score for Llama was significantly lower.
Qualitative analysis: To assess the quality of the features, we use GPT-4o-mini hurst2024gpt as an automated evaluator to analyze the coherence and predictive validity of SAE features from each model. We make sure to ask the LLM to provide explanations justifying each evaluation, a technique shown to improve alignment with human judgments chiang-lee-2023-closer. First, we collected the feature descriptions of the features of the SAE with $\Delta a$ or $\Delta f$ from Neuronpedia [^1] https://www.neuronpedia.org/. Next, we use the prompts in Appendix H, to perform two distinct, complementary evaluations:
@@
Our evaluation reveals significant differences in the SAE features from llama-3.1-8b-it and gemma-29b-it in terms of quality and utility for modeling political contexts. Figure 5 provides an initial, high-level overview of feature quality based on the two key scores: classification confidence and thematic coherence. The "Confidence Score Distribution" (left) shows that the evaluator was able to make high-confidence classifications using features from both models. However, a confidence score of 5 mostly turned out to be where the evaluator could not assign any category, i.e., a confident failure. These “confident failure” outcomes made up 69% for Llama and 39% for Gemma. Beyond this, Gemma’s scores cluster more tightly around 4 and 2 (cases where a category was successfully mapped), while Llama’s distribution has a longer tail toward lower scores, indicating less predictive clarity. The "Feature Coherence Score Distribution" (right) reveals a more dramatic distinction. Llama’s features are consistently rated as having low coherence, with scores tightly clustered around a median of 2. This indicates that its feature sets are frequently perceived as a disjointed mix of unrelated concepts. In contrast, Gemma’s coherence scores are not only higher on average but also show a wider distribution, with an additional concentration of scores between 3.0 and 5.0. This indicates that Gemma is capable of generating highly coherent and interpretable feature sets that represent a unified political or social theme.
Figure 5: Distributions of feature evaluation scores. Gemma’s SAE features achieve higher thematic coherence (right) and stronger predictive validity than Llama’s (left).
Figure 6: TF-IDF–weighted term distributions from SAE feature descriptions. Gemma emphasizes politically and analytically rich terms; Llama focuses on procedural or structural words.
The term frequency plots in Figure 6 provide a quantitative view of the thematic gap between the feature sets. The Llama features are dominated by general, structure-oriented terms such as technical, health, references, and procedural, indicating a focus on syntactic and procedural aspects rather than on semantic or ideological content. Although domain terms like funding and safety appear, they remain embedded in practical rather than interpretive contexts—consistent with lower coherence scores and a descriptive rather than analytical framing. In contrast, Gemma’s features emphasize scientific, political, and analytical discourse, with frequent terms such as political, analysis, data, and discourse highlighting a richer engagement with evaluative and interpretive language. This concentration reflects stronger alignment with socially and politically grounded semantics.
@@
Figure 7: Confusion matrices of topic classification using SAE feature descriptions. Gemma’s features yield clear diagonals and topic discrimination; Llama’s fail to separate categories.
@@
3.4 Ablation of Political Features
To directly test the causal link between political features and the refusal behavior, we performed a zero-ablation of the 71 most salient political features in Gemma-2-9b-it and evaluated their null response rate under the argumentative setting like before. The ablation caused a sharp increase in the model’s refusal rate across most topics, as shown in Figure 8.
Figure 8: Refusal rates before and after ablation of Gemma’s top political features. Removing these features sharply increases refusals, linking capability loss to ideological shallowness.
The ablated model, which originally had a negligible refusal rate, now frequently defaulted to generic safety-aligned responses or failed to follow instructions (though this may be more due to ablation removing some capability, but this was less frequent). This result provides strong causal evidence that refusal can arise from a capability deficit. Removing the specific features that encode political concepts appears to render the model unable to reason about the topic from the instructed perspective, forcing a fallback on its general safety training and mirroring the brittle behavior observed in the feature-poor Llama model.
@@ 4 Discussion
Our findings reveal that ideological depth in LLMs is a measurable structural property reflected in their internal feature representations. Comparing Llama-3.1-8B-IT and Gemma-2-9B-IT, we demonstrate that models exhibit vastly different political feature spaces: Gemma possesses $\sim 7.3\times$ more political features than Llama, directly explaining their behavioral differences. Gemma’s rich feature space enables flexible reasoning across ideological perspectives and high steerability, while Llama’s sparse features result in brittleness and frequent refusals.
Critically, we identify that Llama’s refusal behavior stems from a capability deficit rather than sophisticated safety guardrails. This has important implications for AI alignment. An ideologically shallow model may appear safer through refusal, but this brittleness creates unpredictability – its refusal reflects a failure of understanding, not principled reasoning. Conversely, a feature-rich model like Gemma can more readily adopt undesirable personas, resulting in a more transparent and interpretable behavior, offering better opportunities for monitoring and intervention.
We validate the semantic meaningfulness of these features through automated topic classification using only SAE feature descriptions. The evaluator’s success with Gemma’s features (achieving strong classification performance) confirms they form a coherent political knowledge map, while its failure with Llama’s generic features corroborates our finding of shallow ideological representations.
@@ 5 Conclusion and Future Work
We demonstrate that ideological depth in LLMs is measurable through SAE feature analysis and directly linked to model steerability. Our comparison of Llama-3.1-8B-IT and Gemma-2-9B-IT reveals that Gemma’s larger political feature space enables flexible ideological reasoning, while Llama’s sparse representations result in brittleness and elevated refusal rates. Moreover, we show that targeted ablation of salient political features induces refusal behavior in Gemma, establishing a causal link between feature richness and response capability.
Our future research directions include:
@@
    Mechanistic origins of response instability: Models exhibit significantly higher response variance under conservative personas. Future work should trace the feature circuits responsible for this instability ameisen2025circuit; lindsey2025biology.
@@
    Real-time monitoring: Developing activation probes for political features to enable interpretability and control during inference chen2025personavectorsmonitoringcontrolling.
@@
    Training interventions: Investigating whether ideological depth can be deliberately shaped during fine-tuning or pretraining.
Understanding these latent structures is essential for building LLMs that are simultaneously capable, predictable, and alignable.
6 Limitations
Our analysis relies on Sparse Autoencoders (SAEs), which have known limitations smith2025negative. SAEs heavily depend on training data SAEsAreHighlyDatasetDependent, which could confound our comparisons. However, both SAEs used identical sizes (131K features) and were trained on comparable non-chat corpora (pile-uncopyrighted for GemmaScope, SlimPajama-627B for LlamaScope), providing a largely controlled comparison. The difference in feature magnitude and stark contrast in downstream classification performance suggest genuine representational disparities beyond methodological artifacts. Future work will validate these findings using complementary methods such as probing alain2016understanding, which has shown superior performance on other interpretability tasks.
Additionally, our study focuses on two models of similar size. Whether our findings generalize to larger models or models from different training paradigms remains an open question. It is possible that at a sufficient scale, all models develop rich political feature representations, or conversely, that scaling exacerbates rather than ameliorates representational disparities.
Finally, while we focus on the political domain due to its societal importance and well-defined ground truth labels, ideological reasoning is multifaceted and culturally situated. Our analysis uses US-centric political categories, which may not capture the full complexity of global political discourse. Future work examining non-Western political frameworks would provide a valuable perspective on the generalizability of our findings.
Appendix A Data Curation Process
We used the 1000 prompts from politically-liberal subset of perez2023discovering, consisting of dichotomous questions, intended to identify politically liberal behavior from agentic models. We used a two-step approach to assign a topic for each of the statements: In step 1, we passed all of the statements as a list and prompted Gemini-2.5 pro to group them in categories. However, this results in the LLM regenerating all the statements, which sometimes modifies some of them, a scenario known as context rot hong2025context. Therefore, in step 2, we collect all the categories identified by Gemini and select one statement from the original list, and again prompt the LLM to pick the topic that best matches the statement. To account for selection bias of the LLMs – where they are more likely to pick an option based on their position rather than meaning zheng2024large; rottger2024political – we prompt the LLM 10 times for each statement, each time shuffling the list of categories, and choose the most frequent category for each statement. This ensures that the statements are neatly grouped under distinct topics and no statement is missed or modified during categorization.
After this, all 1000 statements were categorized into 12 different topics.
Appendix B Prompting Conditions
Original persona with no arguments
Original persona with liberal arguments
Original persona with conservative arguments
Liberal persona with no arguments
Liberal persona with liberal arguments
Liberal persona with conservative arguments
Conservative persona with no arguments
Conservative persona with liberal arguments
Conservative persona with conservative arguments
Table 3: Prompting conditions used.
Along with the LLMs’ response to the prompts as their original personas, we also provide detailed instructions to the model to answer the prompts as an individual with either a liberal or a conservative worldview (Appendix G). Additionally, to analyze the ideological depths of the LLMs while answering as different personas, we also evaluate stance consistency in the presence of supporting and counter arguments in the prompt kabir2025wordsreflectbeliefsevaluating. We used Llama-3.3-70B-Instruct to generate at least 5 different supporting and counter arguments for each of the 126 statements in our evaluation set. Combining all these, our approach analyzes 9 conditions using prompt engineering, which are shown in Table 3.
Appendix C Response Consistency
Figure 9: Heatmap of response consistency across nine prompting conditions. Models show the least consistency (most variance) when adopting conservative roles or arguments.
Next, to gauge the political leaning of each model, we evaluated its response coherence measured by Fleiss’ $\kappa$ (Figure 10(a)) and its conservative response rate (Figure 10(b)) for each topic when instructed to respond as a conservative under argumentative pressure.
(a) Response consistency Fleiss’ $\kappa$ over topics.
(b) Conservative response tendency.
Figure 10: Topic-level conservative consistency and tendency under argumentative pressure when instructed to adopt a conservative persona. Gemma achieves stronger conservative alignment but with greater response variance than Llama.
(a) CAA Steering
(b) STA Steering
Figure 11: Topic-level response consistency (Fleiss’ $\kappa$) under activation steering toward the conservative direction ($\text{multiplier}=-1$) with argumentative prompts. Gemma achieves larger conservative shifts but lower consistency.
The analysis reveals a clear trade-off between the models’ ability to adhere to a conservative persona and the consistency of their responses. Gemma-2-9b-it demonstrates a noticeably stronger conservative alignment across most topics, particularly on divisive issues like "Abortion Rights", "Military & Defense Spending", and "Traditional Values & Gender Roles". However, this ability comes at the cost of slightly lower consistency; while it shows significant coherence on "Tax Policy" and "Military & Defense," its consistency drops to fair or poor in several other categories (Figure 10(a)). Conversely, llama-3.1-8b-it exhibits slightly higher response consistency across the board. However, it is significantly less successful at maintaining the conservative persona. It shows its highest conservative tendency on topics like Gun Control and Social Welfare & Poverty but provides predominantly liberal responses on topics like Abortion Rights and Climate & Environment (Figure 10(b)).
We also see this while steering the models to be more conservative using activation steering, where the conservative capability of Gemma still came at the cost of consistency over argumentative pressure. Using Fleiss’ $\kappa$ over the responses at multiplier=1multiplier=-1 with argumentative prompts like before shows this in Figure 11. Llama was better with its response consistency than Gemma on categories like Social Welfare & Poverty and Immigration & Refugees, due to its higher number of liberal or refusal responses on these matters.
Appendix D CAA Layer Sweep
Figure 12: Layer-sweep results for political steerability. Optimal steering layers identified as 14 (Llama) and 20 (Gemma).
For activation steering, we used the well-known MeanDiff method, also known as contrastive activation addition (CAA) rimsky-etal-2024-steering. Using this approach, we sweep over all layers and apply steering with multipliers of -1 and 1. We assess the effect size on the held-out test questions and identify the most effective layer for steering each model’s political ideology. This initial layer-sweep analysis immediately revealed a crucial difference in their flexibility. As shown in Figure 12, Gemma’s probability of a liberal answer could be steered across a wide range (from nearly 100% down to 45%), whereas for Llama it remained confined to a narrow band (96% to 80%). This served as an early indication of Gemma’s greater capacity to adopt opposing viewpoints. Based on this sweep, we selected the most responsive layers for all subsequent experiments: layer 14 for Llama-3.1-8b-it and layer 20 for Gemma-2-9b-it.
Appendix E Factor Analysis of LLM Responses